OpenStudy (anonymous):

is 25+4 radical 7 the correct answer to: (5=2 radical 7)squared if not what is and how did you get it

OpenStudy (anonymous):

when you square rt7 you get 7. the asnwer is 25+10rt7+28 when you foil it out

OpenStudy (anonymous):

so it would be \[53+20\sqrt{7}\] correct??

OpenStudy (anonymous):

53+10rt7

OpenStudy (anonymous):

joojoo, there are two terms of \[10\sqrt{7}\] when you foil it out.

OpenStudy (anonymous):

when you foil it it comes out as this\[25+10\sqrt{7^{2}}+28\]

OpenStudy (anonymous):

the 10 radical 7 is supposed to be all squared not just the 7**

OpenStudy (anonymous):

No it's suppose to be \[20\sqrt{7} \ not 10\sqrt{7}\] So you have \[53+ 20\sqrt{7}\]

OpenStudy (anonymous):

alright thats what i was thinkin after joojoo reminded me it was just the foil method thank you=)

OpenStudy (anonymous):

your welcome